### v1

# Coherence-Mediated Quantum Thermometry in a Hybrid Circuit-QED Architecture

Abstract
Quantum thermometry plays a critical role in the development of low-temperature sensors and quantum information platforms. In this work, we propose and theoretically analyze a hybrid circuit quantum electrodynamics architecture in which a superconducting qubit is dispersively coupled to two distinct bosonic modes: one initialized in a weak coherent state and the other coupled to a thermal environment. We show that the qubit serves as a sensitive readout of the probe mode, mapping the interference between thermal and coherent photon-number fluctuations onto measurable dephasing. This mechanism enables enhanced sensitivity to sub-millikelvin thermal energy fluctuations through Ramsey interferometry. We derive analytic expressions for the qubit coherence envelope, compute the quantum Fisher information for temperature estimation, and demonstrate numerically that the presence of a coherent reference amplifies the qubit’s sensitivity to small changes in thermal photon occupancy. Our results establish a new paradigm for quantum-enhanced thermometry and provide a scalable platform for future calorimetric sensing in high-energy physics and quantum metrology.
I Introduction
Quantum thermometry has emerged as a critical capability in modern quantum science and technology. It plays a central role in the calibration of cryogenic environments, the characterization of quantum devices and sensors, and the detection of weak energy events in low-temperature physics [1, 2, 3, 4]. Beyond practical applications, precise temperature measurement also provides a platform for exploring fundamental thermodynamic limits in the quantum regime. Superconducting quantum circuits, particularly those operating in the microwave domain, offer a versatile and scalable platform for implementing such thermometric protocols [5, 6]. In these systems, qubits serve as exquisitely sensitive probes of their electromagnetic environment and can detect thermal photon populations through their decoherence dynamics [7, 8, 9, 10, 11].
The most common approach to qubit-based thermometry exploits the dispersive interaction between a qubit and a thermalized cavity mode. In the dispersive regime, photon-number fluctuations induce pure dephasing of the qubit, which can be monitored via Ramsey or spin-echo sequences [12, 13]. The decay of qubit coherence directly encodes the thermal occupancy $\bar{n}$ of the mode and thereby its effective temperature. This single-mode dephasing protocol has been demonstrated in a variety of circuit quantum electrodynamics (cQED) architectures [14, 15, 16] and is particularly effective when $\bar{n}\gtrsim 1$. However, in the ultra-low-temperature limit $\bar{n}\ll 1$, corresponding to mode temperatures below $50$ mK at GHz frequencies, the variance of Bose–Einstein photon statistics becomes exponentially suppressed, and the sensitivity of this method is fundamentally limited by the vanishing signal-to-noise ratio of thermal fluctuations.
Here we introduce a two-mode quantum thermometry protocol that overcomes this low-temperature sensitivity floor by leveraging coherent–thermal interference. Specifically, we consider a hybrid cQED architecture in which a superconducting qubit is dispersively coupled to two bosonic modes: a high-$Q$ 3D cavity initialized in a weak coherent state, and a low-$Q$ planar resonator coupled to a thermal environment. The coherent mode acts as a stable phase reference, while the thermal mode functions as the sensing arm. The qubit, coupled to the probe cavity, serves as a phase-sensitive readout that acquires a stochastic phase shift determined by the joint photon-number fluctuations of both modes. Therefore, the qubit’s coherence decay directly reflects the interference between their photon statistics.
This architecture realizes a novel interferometric thermometry scheme in which small thermal signals are amplified by embedding them in the phase dynamics of a coherent reference. We analytically derive the coherence envelope of the probe mode as a function of interaction time, thermal photon number, and coherent amplitude, and show how it is faithfully mapped onto qubit Ramsey dephasing through dispersive readout. The resulting dephasing dynamics depend not only on the variances of each mode individually but also on their interference, enabling thermal signal amplification through coherent–thermal correlations. To quantify the resulting sensitivity, we calculate the quantum Fisher information (QFI) associated with qubit coherence measurements and identify optimal working points that maximize temperature resolution [17, 18, 19, 20].
Numerical simulations based on exact unitary interaction validate the analytic model and show that sub-millikelvin sensitivity is, in principle, achievable with realistic circuit parameters. The proposed architecture is compatible with existing cQED platforms, including hybrid designs that combine planar resonators and qubits with 3D cavities [21, 22, 23], and does not require photon counting or non-Gaussian state preparation. These features make it a scalable and noise-resilient foundation for thermometry at ultra-low energies. Potential applications range from on-chip quantum diagnostics [24] and cryogenic calorimetry [25] to thermal signal amplification in rare-event detection platforms such as axion or hidden-photon dark matter searches [26, 27]. Taken together, these results establish coherent–thermal interferometry as a promising route toward quantum-limited thermometric sensing in superconducting circuits.
II Theory: Thermal-to-Coherent Mapping
Quantum thermometry aims to extract information about temperature from quantum systems with minimal disturbance and maximal sensitivity. In this work, we consider a hybrid bosonic system where a thermal field is entangled with a coherent reference field. This interaction, acting as a noise-to-phase transduction mechanism, enables temperature-dependent fluctuations in the thermal field to be imprinted onto a well-controlled coherent mode, thereby amplifying otherwise weak thermal signals and allowing them to be efficiently detected by monitoring the qubit’s coherence. Below we present a detailed theoretical framework for this thermal-to-coherent mapping mechanism.
II.1 System Model
Figure 1: Schematic of the proposed thermometry architecture. A low-$Q$ thermal resonator $\hat{a}$ is thermalized by a bath at temperature $T$. Through an engineered cross-Kerr interaction $\lambda$, realized with two fixed-frequency transmons ($Q_{1}$, $Q_{2}$), fluctuations of $\hat{a}$ are mapped onto a high-$Q$ 3D cavity probe mode $\hat{b}$. A sensing transmon couples strongly to the probe with dispersive strength $\chi_{b}$ for readout, while any residual coupling $\chi_{a}$ to the thermal mode acts as a parasitic dephasing channel. Temperature information encoded in the probe can be extracted either via qubit Ramsey (coherence-mediated) or direct heterodyne detection (phase-shift).
As shown in Fig. 1, we consider a hybrid system comprising a qubit coupled to two bosonic modes:
Mode $\hat{a}$: a thermal field at temperature $T$, initialized in a thermal state $\rho_{\mathrm{th}}(T)$ with mean photon number $\bar{n}_{a}$. A low-quality 2D resonator, such as a CPW or lumped-element LC resonator, can be used to realize this mode.
Mode $\hat{b}$: a coherent reference field prepared in a coherent state $|\alpha\rangle$ with amplitude $\alpha$. This mode can be implemented using an ultra-high-quality 3D cavity.
The qubit interacts dispersively with both modes. In addition, a nonlinear cross-Kerr interaction mediates a joint coupling between the thermal and coherent modes. The total interaction Hamiltonian is given by
$\hat{H}_{\mathrm{int}}=\hat{\sigma}_{z}\left(\chi_{a}\hat{n}_{a}+\chi_{b}\hat{n}_{b}\right)+\lambda\hat{n}_{a}\hat{n}_{b},$ (1)
where $\hat{n}_{a(b)}=\hat{a}^{\dagger}\hat{a}$ ($\hat{b}^{\dagger}\hat{b}$) are the photon number operators, $\chi_{a}$ and $\chi_{b}$ are dispersive coupling strengths between the qubit and each mode, and $\lambda$ is the cross-Kerr strength. The first two terms represent standard dispersive phase shifts proportional to photon number, while the final term provides a direct photon-number–dependent coupling between the two modes. Here $\lambda$ arises as an effective interaction mediated by virtual transitions through multiple qubit pathways [28, 29, 30]. In this work, we intentionally decouple the thermal field from the qubit so $\chi_{a}$ is negligible. This avoids qubit decoherence due to direct thermal contact, while preserving thermal sensitivity via the indirect $\lambda$-mediated path.
II.2 Lamb-Like Shift: Thermal-to-Coherent Mapping
The cross-Kerr interaction, $\lambda\hat{n}_{a}\hat{n}_{b}$, couples the thermal mode $\hat{a}$ to the coherent probe mode $\hat{b}$, such that thermal photon-number fluctuations induce frequency shifts in the coherent mode [31, 32]. While the traditional Lamb shift arises as a second-order correction to a quantum system’s energy levels due to vacuum fluctuations, the present scenario exhibits a reversed structure: it is the coherent field, rather than the qubit, that experiences environment-induced shifts. Here, the thermal mode acts as an effective fluctuating environment for the probe.
From the perspective of the coherent mode, the effective Hamiltonian reads
$\hat{H}_{b}^{\mathrm{eff}}=\left(\omega_{b}+\lambda\hat{n}_{a}\right)\hat{n}_{b},$ (2)
where the thermal occupation $\hat{n}_{a}$ follows Bose–Einstein statistics: $\bar{n}_{a}(T)=1/[\text{exp}(\hbar\omega_{a}/k_{B}T)-1].$ Due to the probabilistic nature of thermal fluctuations, each realization of $n_{a}$ photons shifts the frequency of the coherent mode by $\lambda n_{a}$, resulting in a distribution of frequency shifts centered at $\omega_{b}+\lambda\bar{n}_{a}$.
This random frequency shift leads to stochastic phase accumulation. For a single realization with $n_{a}$ photons, the phase shift after time $\tau$ is deterministic: $\phi_{b}(\tau,T)=\lambda n_{a}\tau$. Across the thermal ensemble, however, this becomes a random variable: $\phi_{b}=\int_{0}^{\tau}\delta\omega_{b}(t)\,dt,$ where $\delta\omega_{b}(t)$ captures the instantaneous fluctuation in frequency due to the thermal photon number. This process defines a pure dephasing channel for the coherent state: it causes a gradual loss of phase coherence without energy dissipation.
Ensemble-averaging over many thermal realizations transforms the initial pure coherent state into a statistical mixture of phase-rotated states. As a result, the coherent mode no longer exhibits a well-defined global phase but instead acquires a broadened phase distribution. This dephasing is observable as a decay in the coherent amplitude, and it plays a central role in Strategy 1, where we model the coherence envelope as a temperature-sensitive observable.
Importantly, while the original Lamb shift arises as a second-order effect ($\propto\lambda^{2}$), the thermally induced frequency shift in our model is a first-order effect in $\lambda$, but becomes effectively stochastic due to the underlying thermal noise. The resulting dephasing thus encodes temperature-dependent fluctuations into the probe’s coherence dynamics.
II.3 Strategy 1: Coherence-mediated envelope and QFI
After interacting with the thermal mode via a cross-Kerr coupling, the coherent probe acquires temperature-dependent phase noise. This dephased coherent state is interrogated by a dispersively coupled qubit using a Ramsey sequence [33, 34]. The qubit’s coherence reflects the fluctuating probe phase:
$|\psi_{q}(\tau)\rangle=\frac{1}{\sqrt{2}}\left(|0\rangle+e^{i\phi_{b}}|1\rangle\right),$ (3)
where the accumulated phase $\phi_{b}=\lambda n_{a}\tau$ originates from thermal fluctuations and is therefore temperature dependent.
We model the probe mode as undergoing random phase kicks induced by photon-number fluctuations in the thermal mode. Since $\phi_{b}\propto n_{a}$, its statistics are inherited from the Bose–Einstein distribution of $n_{a}$, with mean $\bar{n}_{a}$ and variance $\mathrm{Var}(n_{a})=\bar{n}_{a}(\bar{n}_{a}+1)$. Accordingly, the probe phase has mean $\langle\phi_{b}\rangle=\lambda\tau\bar{n}_{a}$ and variance
$\sigma_{\phi}^{2}=\mathrm{Var}(\phi_{b})=(\lambda\tau)^{2}\bar{n}_{a}(\bar{n}_{a}+1).$ (4)
In the low-temperature limit ($\bar{n}_{a}\ll 1$), the variance reduces to $\sigma_{\phi}^{2}\approx(\lambda\tau)^{2}\bar{n}_{a}$.
Although $n_{a}$ is not Gaussian distributed, the accumulated phase noise results from many small, independent fluctuations and can be approximated by Gaussian diffusion (central limit theorem) [1, 35]. The probe is thus described as a phase-averaged mixture,
$\rho_{b}(T)=\int d\phi\,\frac{e^{-\phi^{2}/2\sigma_{\phi}^{2}}}{\sqrt{2\pi\sigma_{\phi}^{2}}}\,|\alpha e^{i\phi}\rangle\langle\alpha e^{i\phi}|,$ (5)
which captures the temperature-dependent dephasing of the probe field. This Gaussian phase-diffusion picture is directly analogous to qubit dephasing from a dispersively coupled thermal cavity, with coherence loss governed by the same $\bar{n}_{a}(\bar{n}_{a}+1)$ scaling [35, 36, 37, 38].
Figure 2: Simulated coherence-based QFI $\mathcal{F}_{C}$ and corresponding temperature sensitivity $\delta T$, with thermal frequency $\omega_{a}/2\pi=1~\mathrm{GHz}$, coherent amplitude $\alpha=2.0$, and cross-Kerr coupling $\lambda/2\pi=50~\mathrm{kHz}$. (a) Heatmap of $\mathcal{F}_{C}$ as a function of thermal bath temperature $T$ and interaction time $\tau$. The red dashed line indicates the optimal $\tau(T)$ that maximizes $\mathcal{F}_{C}$ for each temperature. (b) Minimum detectable temperature change $\delta T_{\mathrm{min}}=1/\sqrt{\nu\mathcal{F}_{C}}$ (blue solid line) as a function of temperature, computed at the optimal $\tau$ (green dashed line). A sensitivity of approximately $60~\mu\mathrm{K}$ is achieved near $T\sim 10~\mathrm{mK}$ with $\tau=10~\mu\mathrm{s}$. In this simulation, we set the number of measurement repetitions to $\nu=10^{4}$.
The probe coherence envelope under Gaussian phase diffusion is
$\mathcal{V}_{C}(\tau,T)\equiv C(\tau,T)=\exp\left[-2|\alpha|^{2}\left(1-e^{-\Gamma_{\phi}}\right)\right],$ (6)
with the effective dephasing rate $\Gamma_{\phi}=\sigma_{\phi}^{2}\approx(\lambda\tau)^{2}\bar{n}_{a}$ (Appendix B).
This expression reflects the progressive decoherence of a coherent state subject to Gaussian phase fluctuations with temperature-dependent variance, and it is experimentally accessible via Ramsey measurements of the qubit. At low temperatures ($\bar{n}_{a}\to 0$), dephasing vanishes and $C(\tau,T)\to 1$. At high temperatures, the envelope saturates at $\exp(-2\alpha^{2})$, reflecting complete phase scrambling [12]. The nonlinear dependence on both $\alpha$ and $\bar{n}_{a}$ enables strong sensitivity to thermal fluctuations even in the sub-photon regime.
To quantify thermometric performance, we compute the quantum Fisher information (QFI) associated with the measurable signal $C$: $\mathcal{F}_{C}(T)=|\partial_{T}C|^{2}/(1-C^{2}),$ with $\Phi(T)$ constant (Appendix A). Substituting $C(\tau,T)$ gives
$\mathcal{F}_{C}(T)=\frac{\left[2\alpha^{2}e^{-\Gamma_{\phi}}(\lambda\tau)^{2}\,\partial_{T}\bar{n}_{a}\right]^{2}C^{2}}{1-C^{2}},$ (7)
where $\partial_{T}\bar{n}_{a}=\hbar\omega_{a}/\!\left[\bar{n}_{a}(\bar{n}_{a}+1)k_{B}T^{2}\right]$. Eq. 7 establishes the fundamental precision bound of the thermometer [17, 19, 20, 18, 39].
Physically, the probe converts temperature-dependent photon-number fluctuations into measurable phase diffusion. In this coherence-mediated model, the QFI increases with $\alpha^{2}$ but is simultaneously limited by visibility decay in the denominator. In the weak-dephasing regime ($\Gamma_{\phi}\ll 1$), one finds $\mathcal{F}_{C}\;\approx\;\alpha^{2}(\lambda\tau)^{2}\,\bar{n}_{a}\left(\hbar\omega_{a}/k_{B}T^{2}\right)^{2}.$ Thus, the coherent-probe scheme provides a tunable transduction “gain” via $\alpha$ (and $\tau$), but this gain is bounded in practice by visibility decay, available probe power, bandwidth, and backaction. We therefore interpret the improvement as a resource-assisted sensitivity, which is achieved by allocating probe photons to strengthen the temperature-to-phase conversion, rather than as an unlimited amplification.
Figure 2 shows a simulated heatmap of $\mathcal{F}_{C}$, highlighting the optimal interaction time $\tau(T)$ shown as dashed line in panel(a). Temperature sensitivity is defined as $\delta T_{\mathrm{min}}=1/\sqrt{\nu\mathcal{F}_{C}(T)},$ with $\nu$ the number of independent repetitions. For realistic parameters ($\lambda/2\pi=50~\mathrm{kHz}$, $\omega_{a}/2\pi=1~\mathrm{GHz}$, and $\alpha=2$), the thermometer achieves a sensitivity $\delta T\sim 60~\mu\mathrm{K}$ near $T=10~\mathrm{mK}$, with an optimal interaction time around $\tau\sim 10~\mu\mathrm{s}$ and measurement repetition $\nu=10^{4}$. Such timescales are readily supported by high-Q 3D cavities, which exhibit photon lifetimes up to tens of milliseconds [40, 41, 23].
II.4 Strategy 2: Probe Phase-Shift Tracking
[…]
Building on the qubit-based coherence measurement in Strategy 1, the same temperature-dependent phase $\phi_{b}(T)=\lambda\tau\bar{n}_{a}(T)$ can be viewed as the parameter that would accumulate during a Ramsey sequence on a dispersively coupled qubit. Alternatively, we measure the phase variation directly via heterodyne detection of the high-$Q$ probe cavity, instead of mapping it onto a qubit superposition.
In this heterodyne-based implementation, the cross-Kerr interaction between the thermal mode and probe induces a temperature-dependent phase shift of the intracavity field. This phase is faithfully transferred to the outgoing field and measured through heterodyne detection of its quadratures. Because the amplitude remains unchanged, the probe visibility is unity,
$\mathcal{V}_{\Phi}(\tau,T)=1,$ (8)
and the information about temperature resides purely in the phase. The corresponding quantum Fisher information,
$\mathcal{F}_{\Phi}(T)=\big(\lambda\tau\,\partial_{T}\bar{n}_{a}\big)^{2},$ (9)
represents the ultimate precision bound for phase-based thermometry (Appendix C). Figure 3 shows $\mathcal{F}_{\Phi}(T)$ and the resulting temperature sensitivity $\delta T(T)=1/\sqrt{\nu\mathcal{F}_{\Phi}(T)}$ for several $\tau$. At low $T$, $\mathcal{F}_{\Phi}$ is exponentially suppressed, $\mathcal{F}_{\Phi}\!\sim\!T^{-4}e^{-2\hbar\omega_{a}/k_{B}T}$, while at high $T$ it saturates to $(\lambda\tau k_{B}/\hbar\omega_{a})^{2}$. Longer interaction times simply rescale the vertical axis as $\mathcal{F}_{\Phi}\propto\tau^{2}$ and $\delta T\propto 1/\tau$.
This approach leverages the exceptional phase stability of ultra-high-$Q$ 3D cavities ($\kappa_{b}\!\ll\!$ kHz) to permit millisecond-scale integration times without qubit-coherence limitations. Thus, it naturally supports continuous, low-bandwidth monitoring of slow thermodynamic drifts with state-of-the-art sensitivity.
Beyond steady-state thermometry, the same phase-tracking technique enables quantum calorimetry [42]. A discrete energy deposition in the absorber produces a sudden jump $\Delta n_{a}$ in the thermal occupation, resulting in an instantaneous probe-phase step $\Delta\phi_{b}=\lambda\tau\Delta n_{a}$. Continuous heterodyne monitoring allows such steps to be time-tagged and resolved above the noise floor, enabling detection of rare, quantized energy arrivals. This connects our architecture to calorimetric applications in nuclear and high-energy physics [43, 44], where resolving single-particle energy deposits with minimal backaction is essential. The long averaging time $\tau$ that optimizes steady-state sensitivity is, in the calorimetric mode, replaced by an effective integration window set by the probe bandwidth and digital filter, which together determine both energy resolution and timing jitter.
Taken together, the phase-shift strategy complements the coherence-mediated approach. While Strategy 1 enables fast, high-bandwidth thermometry with short interaction times, Strategy 2 exploits long-lived cavity coherence for ultra-high sensitivity and naturally extends to calorimetry, where discrete energy events are registered as probe phase jumps.
III Experimental Implementation and Feasibility
Building on the theoretical framework outlined above, we now describe a realistic experimental implementation of the coherence-mediated quantum thermometry scheme.
III.1 Circuit Architecture and Mode Realization
The proposed thermometer comprises three essential elements: a high-coherence superconducting qubit, a low-frequency thermal mode, and a high-frequency coherent probe mode. These are integrated in a hybrid architecture combining planar CPW resonators with ultra-high-$Q$ 3D cavity components.
The probe is implemented as a 3D superconducting cavity in the 6106{-}10 GHz band with photon lifetimes in milli-second range [40, 41, 23]. Such longevity enables interaction times far exceeding qubit coherence times, thereby enhancing thermometric sensitivity. The long-lived probe state reduces the need for frequent reinitialization and allows repeated qubit interactions with minimal backaction. Compared to planar resonators, 3D cavities exhibit negligible internal dissipation, suppressing excess noise. Their dispersive coupling strength $\chi_{b}$ can be engineered geometrically (via qubit placement in the field) or spectrally (via detuning). They are also compatible with high-fidelity microwave readout, supporting both direct heterodyne detection and qubit-mediated measurements.
The thermal channel is realized as a deliberately lossy CPW resonator or lumped LC circuit at 121{-}2 GHz. Operating at lower frequencies ensures a sizable Bose occupation $\bar{n}_{a}$ even at tens of millikelvin. By coupling strongly to an engineered bath, such as a terminated transmission line or resistor load, the mode rapidly equilibrates with its environment [24, 45], making it a faithful temperature sensor.
[…]
In the coherence-mediated scheme, the qubit remains in its ground state while thermal fluctuations imprint dephasing noise onto the long-lived probe during an interaction interval $\tau$. A subsequent Ramsey-type sequence maps this probe dephasing back to the qubit for readout. In the phase-shift scheme, the probe field itself is monitored continuously via heterodyne detection; its ultra-high $Q$ enhances sensitivity to stochastic fluctuations in the thermal mode occupancy $\bar{n}_{a}(T)$.
III.2 Feasibility: Cross-Kerr Coupling and Practical Constraints
In practice, feasibility is set by the need to generate measurable probe responses within realistic interaction times, while staying compatible with coherence times, cavity linewidths, and achievable qubit–cavity detunings in state-of-the-art superconducting platforms. These constraints define the operating window where both cross–Kerr couplings and dispersive pulls can be harnessed for thermometry in a competitive regime.
III.2.1 Engineering Enhanced Cross–Kerr
[…]
In our scheme, the effective cross–Kerr must be strong enough ($\lambda/2\pi\sim 10$$50~\mathrm{kHz}$) to imprint temperature–dependent fluctuations onto the probe within a realistic interaction time $\tau$, while remaining compatible with achievable detunings and coherence in a hybrid circuit–QED architecture.
Conventional approaches use four–wave mixing to enhance the interaction between modes $a$ and $b$. An off–resonant pump can dress an existing transmon and increase the cross–Kerr without new hardware [31, 48]. Additionally, driven couplers placed between the modes can offer comparable, tunable $\lambda$ [49, 50, 51]. While flexible, these methods introduce AC–Stark and Kerr drifts and can raise the inverse–Purcell/noise floor seen by a high–$Q$ probe, adding calibration and packaging overhead that is difficult to reconcile with a 3D–cavity implementation.
We therefore implement $\lambda\,\hat{n}_{a}\hat{n}_{b}$ using two fixed–frequency transmons (Q1, Q2) that passively bridge the thermal mode $a$ and the probe $b$ (see Fig. 1). Mode $a$ couples dispersively to Q1 with pull $\chi_{a1}=g_{a1}^{2}/\Delta_{a1}$, mode $b$ couples to Q2 with pull $\chi_{b2}=g_{b2}^{2}/\Delta_{b2}$, and the qubits are capacitively linked with an XY exchange $J_{XY}$. In the deep–dispersive regime, $|g_{a1,b2}|\ll|\Delta_{a1,b2}|,\quad|J_{XY}|\ll|\Delta_{12}|,$ the virtual coupling chain $a-Q_{1}-Q_{2}-b$ yields an effective cross–Kerr (Appendix E)
$\lambda\;=\;\frac{8\,\chi_{a1}\,\chi_{b2}\,J_{XY}^{2}}{\Delta_{12}^{3}}\;\times\;\big[1+\mathcal{O}(\tfrac{\Delta_{12}}{\Delta_{a1}},\tfrac{\Delta_{12}}{\Delta_{b2}})\big],$ (10)
where the order–unity corrections reflect finite transmon anharmonicities and detuning signs [31, 52]. Figure 4 shows the effective cross-Kerr $\lambda$ as a function of dispersive pulls $\chi_{a1,b2}$. For a fixed dispersive ratio $\rho\equiv J_{XY}/\Delta_{12}$, the scaling rewrites as $\lambda=8\,\chi_{a1}\chi_{b2}\rho^{2}/\Delta_{12}$, so increasing $\chi_{a1}$ (thermal side) and/or reducing $\Delta_{12}$ most efficiently boosts $\lambda$ while keeping $\chi_{b2}$ modest to protect the probe $Q$.
[…]
III.2.2 Impact of Qubit–Mode Couplings on Visibility
Figure 5: Qubit Ramsey visibility as a function of dispersive coupling strengths during the Ramsey window $\tau_{R}$. (a) Dependence on $\chi_{b}/2\pi$ for 10, 20, 5010,\,20,\,50 kHz. (b) Dependence on $\chi_{a}/2\pi$ for 1, 5, 101,\,5,\,10 kHz. The y-axis shows the normalized Ramsey visibility (qubit contrast) versus the Ramsey integration time $\tau_{R}$. Larger dispersive couplings increase dephasing during readout, shortening the usable $\tau_{R}$ window.
In the coherence-mediated scheme, the qubit is prepared and held in its ground state throughout the imprinting interval $\tau$, so it does not acquire dephasing from either mode during sensing. Visibility becomes relevant only during the subsequent Ramsey readout of duration $\tau_{R}$, when the qubit is placed in a superposition to sample the probe’s phase fluctuations.
If the qubit couples dispersively to a bosonic mode $m\in\{a,b\}$ with photon-number variance $\mathrm{Var}(n_{m})$ and linewidth $\kappa_{m}$, a second-order cumulant expansion gives the coherence envelope during the Ramsey window (Appendix F)
$C_{q}^{(m)}(\tau_{R})=\exp\!\left\{-\,4\chi_{m}^{2}\,\mathrm{Var}(n_{m})\left[\frac{\kappa_{m}\tau_{R}-1+e^{-\kappa_{m}\tau_{R}}}{\kappa_{m}^{2}}\right]\right\},$
with $\mathrm{Var}(n_{b})=|\alpha|^{2}+\bar{n}_{b}(1+\bar{n}_{b})$ for the probe and $\mathrm{Var}(n_{a})=\bar{n}_{a}(1+\bar{n}_{a})$ for the thermal mode. The coherent amplitude $|\alpha|^{2}$ contributes photon shot noise.
[…]
In summary, no dynamical decoupling is required during sensing because the qubit stays in $|g\rangle$ throughout $\tau$. The only exposure to dephasing occurs in the short Ramsey interval $\tau_{R}$, where a large $\chi_{b}$ enables fast, high-fidelity readout, while a small $\chi_{a}$ avoids parasitic dephasing.
IV Comparative Performance of Quantum Thermometry Strategies
Figure 6: Comparison of thermometric strategies. (a) Quantum Fisher information (QFI) $\mathcal{F}$ as a function of interaction time $\tau$ for the coherence-mediated (blue), phase-shift (orange), and qubit-only Ramsey (green) schemes. (b) Normalized coherence envelopes. The blue curve shows the probe envelope $C(\tau,T)$ in the coherence-mediated scheme, which is mapped to the qubit through dispersive readout. The green curve shows the qubit’s direct Ramsey envelope under thermal dephasing. The phase-shift scheme maintains unit contrast (orange). Together, the plots highlight that the coherence-mediated scheme achieves higher QFI at short times, while visibility decay ultimately limits long-$\tau$ sensitivity.
To benchmark performance, we compare the QFI and the associated visibility across three strategies: coherence-mediated sensing, direct phase-shift tracking, and qubit-only Ramsey thermometry [7, 8]. Details of the qubit-only scheme are provided in Appendix D. Figure 6 summarizes the results.
Panel (a) shows the QFI $\mathcal{F}_{T}(\tau)$ at $T=10$ mK. The coherence-mediated strategy (blue) exhibits the largest QFI at short and intermediate times, owing to the probe’s ability to convert thermal number fluctuations into amplified phase diffusion. However, this gain collapses at long $\tau$ as the probe visibility decays. The phase-shift strategy (orange) accumulates information linearly with $\tau$ and avoids dephasing, yielding steadily increasing QFI but with smaller values at short times. The qubit-only scheme (green) lies in between: it achieves an intermediate maximum QFI that can surpass the phase-shift strategy at moderate $\tau$, but is ultimately limited by qubit dephasing from the thermal mode.
Panel (b) displays the corresponding visibilities. The coherence-mediated scheme suffers the fastest decay due to probe dephasing. The qubit-only scheme also loses visibility, though more gradually. By contrast, the phase-shift scheme maintains unit visibility at all times, since the signal resides entirely in deterministic phase accumulation.
Taken together, these results highlight complementary trade-offs among the three strategies. The coherence-mediated scheme uniquely exploits the qubit’s quantum coherence as a transient amplifier, transducing thermal fluctuations into enhanced phase sensitivity. Within the weak-dephasing window ($\tau\lesssim 10~\mu\mathrm{s}$ for the chosen parameters, extendable by reducing the cross-Kerr $\lambda$), this approach achieves the strongest instantaneous quantum Fisher information and the steepest signal slope, outperforming both qubit-only and phase-shift tracking schemes in the short-time regime. In contrast, the qubit-only method is fundamentally limited by intrinsic dephasing, leading to a comparable visibility–sensitivity trade-off at longer $\tau$. The phase-shift strategy is the most robust, supporting arbitrarily long interaction times, though it lacks the amplification benefit of a coherent probe. Thus, coherence-mediated thermometry provides the greatest quantum gain, contingent on a high-$Q$ probe and precise dephasing control, both of which are well within current experimental reach, making it a promising route for practical quantum thermometry.
The essential difference between these two schemes lies in how temperature information is stored. In the qubit-only approach, the qubit is both sensor and memory: the same thermal-induced dephasing that encodes $T$ simultaneously destroys qubit visibility as $\tau$ grows. Because this dephasing originates from the sensing channel itself, dynamical decoupling cannot mitigate it without erasing the signal. By contrast, in the coherence-mediated scheme, the qubit can remain isolated during the sensing interval while thermal fluctuations are mapped onto the probe. The qubit is only re-engaged briefly for readout, protected by dynamical decoupling or idling during interaction. A high-$Q$ probe cavity suppresses visibility loss, enabling longer interaction times before decoherence dominates. In this way, the coherence-mediated strategy shifts the fundamental limitation away from the qubit’s intrinsic coherence and into a regime where cavity quality factors and readout protocols can be engineered for improved sensitivity.
Beyond raw sensitivity, a practical figure of merit is the information gained per unit wall-clock time. We define the Fisher-information rate as $\dot{\mathcal{F}}_{T}(\tau)\equiv\mathcal{F}_{T}(\tau)/(\tau+\tau_{\mathrm{oh}}),$ where $\tau_{\mathrm{oh}}$ accounts for preparation and readout overhead.
For the coherence-mediated and qubit-only schemes, the scaling $\mathcal{F}_{T}\!\sim\!\tau^{2}e^{-2\Gamma\tau}$ yields an optimal interaction time $\tau^{\star}\!\sim\!1/(2\Gamma)$ that maximizes $\dot{\mathcal{F}}_{T}$ and defines an effective measurement bandwidth $B^{\star}\!\sim\!\Gamma$. In contrast, the phase-shift strategy, with $\mathcal{F}_{T}\!\sim\!\tau^{2}$ and no visibility penalty, supports a continuously increasing $\dot{\mathcal{F}}_{T}(\tau)$ up to the limit set by technical drifts. Rather than a bandwidth constraint, this reflects its ability to integrate temperature information coherently over long durations, making it ideally suited for quantum calorimetry where slow thermal relaxation or discrete energy-deposition events are to be resolved with maximal precision. Thus, the phase-tracking scheme complements the high-bandwidth coherence-mediated approach by providing a low-noise, high-fidelity pathway toward long-term thermodynamic monitoring.
This comparison shows that the coherence-mediated and qubit-only protocols naturally operate in a high-bandwidth regime, making them well suited for fast thermometry and real-time temperature tracking. By contrast, the phase-shift strategy excels in the opposite limit: it enables coherent accumulation of temperature information over long durations without visibility loss, providing the optimal platform for quantum calorimetry and other applications requiring ultra-stable, long-integration measurements.
V Outlook and Conclusion
The principle demonstrated here, using a coherent ancillary mode as an information buffer between a noisy environment and a fragile qubit, defines a general paradigm for quantum-enhanced thermometry and noise spectroscopy. In practice, the scheme is compatible with state-of-the-art circuit-QED hardware, where high-$Q$ 3D cavities routinely provide interaction times of tens of milliseconds and cross-Kerr couplings in the 10105050 kHz range can be engineered with multi-qubit pathways. Beyond temperature estimation, related strategies may be adapted to quantum calorimetry, fluctuation spectroscopy, or the characterization of correlated noise in quantum processors. More broadly, coherence-mediated sensing architectures could provide useful tools in nuclear and high-energy physics experiments, where ultrasensitive calorimetry at millikelvin scales is urgently needed. Taken together, these considerations point to broad opportunities for coherence-mediated sensing, while also underscoring the immediate impact of our results for quantum thermometry.
In conclusion, we have introduced and analyzed a coherence-mediated thermometry scheme in a hybrid cQED platform, where a thermal mode is monitored indirectly through its cross-Kerr coupling to a coherent probe and read out by a dispersively coupled qubit. By separating information acquisition from readout, the scheme circumvents the intrinsic dephasing bottleneck of qubit-only thermometers, enabling longer interaction times, higher sensitivity, and greater robustness. These results establish qubit-mediated cross-Kerr thermometry as a fundamentally distinct and scalable approach, providing a path toward nondissipative, reusable quantum thermometers with broad applications across quantum science and technology.
Acknowledgements.
This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract no. DE-AC02-07CH11359.
Appendix A QFI from a Coherence Envelope
We consider a qubit probe initialized in the superposition state $|+\rangle=(|0\rangle+|1\rangle)/\sqrt{2}$, which is maximally sensitive to phase interaction. After an interaction time $\tau$, the qubit undergoes pure dephasing due to temperature-dependent fluctuations, leading to the mixed state
$\rho(T)=\frac{1}{2}\begin{pmatrix}1&C(T)\,e^{i\Phi(T)}\\ C(T)\,e^{-i\Phi(T)}&1\end{pmatrix}$, (11)
where $C(T)$ denotes the coherence envelope, i.e. the visibility of Ramsey fringes, and $\Phi(T)$ is a mean accumulated phase. Both quantities generally depend on the interaction time $\tau$; here we suppress this dependence for brevity.
Physically, $C(T)$ captures the loss of coherence induced by temperature-dependent fluctuations (e.g. photon-number noise in a coupled mode), while $\Phi(T)$ represents a deterministic phase shift that can also carry temperature information. In this sense, $C(T)$ quantifies the amplitude of the Bloch vector and $\Phi(T)$ its orientation in the equatorial plane of the Bloch sphere.
The corresponding Bloch vector is
$\mathbf{r}(T)=\big(C\cos\Phi,\,C\sin\Phi,\,0\big),\qquad r=|\mathbf{r}|=C.$
For a general qubit state $\rho=\tfrac{1}{2}(\mathbb{I}+\mathbf{r}\cdot\bm{\sigma})$, the quantum Fisher information (QFI) for estimating a parameter $T$ is known to be [19, 20]
$\mathcal{F}_{T}=\frac{\big(\partial_{T}r\big)^{2}}{1-r^{2}}+r^{2}\big(\partial_{T}\phi\big)^{2},\quad\text{with}\;r=C,\;\phi=\Phi.$ (12)
This expression separates the two distinct resources available for thermometry: changes in the coherence amplitude $C(T)$ and changes in the phase angle $\Phi(T)$.
Explicitly,
$\mathcal{F}_{T}=\frac{\big(\partial_{T}C\big)^{2}}{1-C^{2}}+C^{2}\big(\partial_{T}\Phi\big)^{2}.$ (13)
The first term arises because the eigenvalues of $\rho(T)$, $\lambda_{\pm}=\tfrac{1}{2}(1\pm C)$, depend on $T$ through $C(T)$. The second term originates from the $T$-dependence of the eigenvectors, which rotate with $\Phi(T)$. Thus, loss of coherence (shrinking Bloch vector) and coherent phase shifts (rotation) both provide temperature sensitivity, but they contribute in qualitatively different ways.
In particular, if the accumulated phase $\Phi$ is independent of $T$, the QFI reduces to a purely dephasing-based form:
$\mathcal{F}_{T}=\dfrac{\big(\partial_{T}C\big)^{2}}{1-C^{2}}.$ (14)
This regime describes thermometry strategies where temperature information is encoded solely in the coherence envelope of the qubit.
Although the above derivation was formulated for a qubit undergoing temperature-dependent dephasing, the resulting expression is fully general for any coherence channel characterized by a complex visibility $C(T)e^{i\Phi(T)}$. In the main text, this formalism is applied to the probe mode: the absorber’s temperature modifies the probe’s complex amplitude in the same way, with $C(T)$ describing the loss of visibility and $\Phi(T)$ the temperature-dependent phase shift. The qubit in that scheme serves only as a phase-sensitive transducer that converts the probe’s coherence into a measurable Ramsey signal. Therefore, Eq. (13) quantifies the fundamental temperature information accessible in both the qubit- and probe-based representations.
Appendix B Derivation of the Coherence-Mediated Envelope and QFI
To model the temperature-dependent coherence decay observed in our qubit, we derive the expression for the coherence envelope $C(\tau,T)$ starting from a microscopic picture of phase diffusion induced by thermal fluctuations acting on a coherent state.
We consider a coherent probe mode initially prepared in the state $|\alpha\rangle$, coupled via a cross-Kerr interaction to a thermal mode:
$H_{\text{int}}=\lambda\hat{n}_{a}\hat{n}_{b},$ (15)
where $\lambda$ is the cross-Kerr rate, and $\hat{n}_{a}$, $\hat{n}_{b}$ are photon-number operators for the thermal and coherent modes, respectively.
During an interaction of duration $\tau$, the coherent state accumulates a phase shift depending on the instantaneous thermal photon number:
$|\alpha\rangle\;\longrightarrow\;|\alpha e^{i\phi_{b}}\rangle,\qquad\phi_{b}=\lambda n_{a}\tau.$
Because $n_{a}$ fluctuates in a thermal state, the coherent probe undergoes random phase shifts drawn from a distribution set by the thermal statistics.
For a thermal state, $n_{a}$ follows a Bose–Einstein distribution with mean $\bar{n}_{a}(T)$ and variance $\mathrm{Var}(n_{a})=\bar{n}_{a}\big(1+\bar{n}_{a}\big)$. The resulting phase variance is
$\sigma_{\phi}^{2}=\mathrm{Var}(\phi)=(\lambda\tau)^{2}\bar{n}_{a}(T)\big[1+\bar{n}_{a}(T)\big].$ (16)
Thus, the probe mode experiences Gaussian phase diffusion, described by a phase-averaged state:
$\rho_{b}(T)=\int\!d\phi\,P(\phi)\,|\alpha e^{i\phi}\rangle\langle\alpha e^{i\phi}|,$ (17)
with $P(\phi)=\tfrac{1}{\sqrt{2\pi\sigma_{\phi}^{2}}}\,e^{-\phi^{2}/2\sigma_{\phi}^{2}}.$
The qubit coherence envelope is proportional to the overlap of the original probe state with the diffused mixture:
$C(\tau,T)=\langle\alpha|\rho_{b}(T)|\alpha\rangle=\int d\phi\,P(\phi)\,\big|\langle\alpha|\alpha e^{i\phi}\rangle\big|^{2}.$ (18)
Using the coherent-state overlap $\langle\alpha|\beta\rangle=\exp\!\left[\,\alpha^{*}\beta-\tfrac{1}{2}\big(|\alpha|^{2}+|\beta|^{2}\big)\right],$ with $\beta=\alpha e^{i\phi}$, we obtain
$|\langle\alpha|\alpha e^{i\phi}\rangle|^{2}=\exp\!\big[-2|\alpha|^{2}(1-\cos\phi)\big]=e^{-2\alpha^{2}}e^{2\alpha^{2}\cos\phi}.$
Substituting back, we have
$C(\tau,T)=e^{-2\alpha^{2}}\int d\phi\,P(\phi)\,e^{2\alpha^{2}\cos\phi}.$ (19)
The integral in Eq. (19) can be evaluated in closed form using the Gaussian statistics of $\phi$:
$C(\tau,T)=\exp\!\left[-2\alpha^{2}\left(1-e^{-\sigma_{\phi}^{2}}\right)\right].$ (20)
This expression identifies the effective dephasing strength
$\Gamma_{\phi}(T)=\sigma_{\phi}^{2}=(\lambda\tau)^{2}\bar{n}_{a}(T)\big[1+\bar{n}_{a}(T)\big],$ (21)
which reduces to $\Gamma_{\phi}(T)\approx(\lambda\tau)^{2}\bar{n}_{a}(T)$ for $\bar{n}_{a}\ll 1$. The final coherence envelope is therefore
$C(\tau,T)=\exp\!\left[-2\alpha^{2}\Big(1-e^{-\Gamma_{\phi}(T)}\Big)\right].$ (22)
At low temperatures $\bar{n}_{a}(T)\to 0$], the dephasing vanishes and $C(\tau,T)\to 1$. At high temperatures, the envelope saturates to $\exp(-2\alpha^{2})$, corresponding to complete phase scrambling. The nonlinear dependence on both $\alpha$ and $\bar{n}_{a}$ makes the scheme sensitive to thermal fluctuations even deep in the sub-photon regime. Importantly, unlike the qubit-only protocol where dephasing occurs directly on the qubit, here the qubit can remain dynamically decoupled while the probe collects thermal information. This distinction underlies the extended interaction times achievable in the coherence-mediated scheme when the probe is implemented as a high-$Q$ cavity.
In the coherence-mediated protocol, all $T$-dependence arises from the envelope $C(\tau,T)$; there is no temperature-dependent phase offset $\Phi(T)$. Differentiating Eq. (22) gives
$\partial_{T}C=C\cdot 2\alpha^{2}e^{-\Gamma_{\phi}(T)}\,\Gamma_{\phi}^{\prime}(T),$ (23)
with
$\Gamma_{\phi}^{\prime}(T)=(\lambda\tau)^{2}\partial_{T}\bar{n}_{a},\qquad\partial_{T}\bar{n}_{a}=\frac{\hbar\omega_{a}}{k_{B}T^{2}}\,\bar{n}_{a}(T)\big[1+\bar{n}_{a}(T)\big].$
Substituting in Eq. (14), we obtain
$\mathcal{F}_{T}(\tau,T)=\frac{\Big[2\alpha^{2}e^{-\Gamma_{\phi}(T)}(\lambda\tau)^{2}\,\partial_{T}\bar{n}_{a}\Big]^{2}\,C(\tau,T)^{2}}{1-C(\tau,T)^{2}}.$ (24)
In the low-occupation limit $\bar{n}_{a}\ll 1$ (so $\Gamma_{\phi}\ll 1$ and $C\simeq e^{-2\alpha^{2}(\lambda\tau)^{2}\bar{n}_{a}}$), this simplifies to
$\mathcal{F}_{T}(\tau,T)\;\approx\;\alpha^{2}(\lambda\tau)^{2}\left(\frac{\hbar\omega_{a}}{k_{B}T^{2}}\right)^{2}\bar{n}_{a}(T).$ (25)
This shows that the QFI scales linearly with thermal occupation $\bar{n}_{a}(T)$, quadratically with the effective interaction strength $\lambda\tau$, and linearly with the probe photon number $|\alpha|^{2}$, highlighting the bosonic amplification inherent in the coherence-mediated approach.
Appendix C Probe Phase-Shift QFI
During the interaction time $\tau$, thermal fluctuations in mode $a$ induce a temperature-dependent frequency (and hence phase) shift on the probe mode $b$ via the cross-Kerr coupling $\lambda$. The outgoing probe field, monitored by a heterodyne detector, acquires a mean phase
$\phi_{b}(T)=\lambda\tau\,\bar{n}_{a}(T),$ (26)
where $\bar{n}_{a}(T)$ is the thermal occupation of the absorber mode. The measurement outcome is the complex probe quadrature $b_{\text{out}}=|\alpha|e^{i\phi_{b}(T)}$, whose phase carries the temperature information.
The temperature-dependent phase shift transforms an initial coherent probe $|\alpha\rangle$ into
$|\psi(T)\rangle=|\alpha e^{i\phi_{b}(T)}\rangle.$ (27)
The quantum Fisher information (QFI) for temperature estimation from this pure state is [17, 53, 54]
$\mathcal{F}_{\Phi}(T)=4\!\left[\langle\partial_{T}\psi|\partial_{T}\psi\rangle-\big|\langle\psi|\partial_{T}\psi\rangle\big|^{2}\right].$ (28)
Using $\partial_{T}|\psi\rangle=i(\partial_{T}\phi_{b})\,\hat{n}_{b}|\psi\rangle$ and the photon-number variance $\mathrm{Var}(\hat{n}_{b})=|\alpha|^{2}$ for a coherent state, we find
$\mathcal{F}_{\Phi}(T)=4|\alpha|^{2}\big(\partial_{T}\phi_{b}(T)\big)^{2}.$ (29)
Thus, the thermometric sensitivity of the heterodyne phase-tracking scheme grows quadratically with the temperature derivative of the induced phase. Although increasing intracavity photons $|\alpha|^{2}$ enhances $\mathcal{F}_{\Phi}$ linearly, it introduces potential backaction and power-induced dephasing. To remain in the weak-measurement regime, we set $|\alpha|^{2}=1$ in the following expressions. We also choose to absorb the factor of $4$ into the operational sensitivity definition used for comparison across schemes.
Finally, inserting the explicit dependence of $\phi_{b}(T)=\lambda\tau\bar{n}_{a}$ gives the compact form
$\mathcal{F}_{\Phi}(T)=\big(\lambda\tau\,\partial_{T}\bar{n}_{a}\big)^{2},$ (30)
which quantifies the thermometric sensitivity of the phase-shift scheme in the single-photon normalization.
Appendix D Derivation of Qubit-Only-Based Coherence Envelope and QFI
[…]
We consider a qubit dispersively coupled to a single harmonic mode $\hat{a}$ of frequency $\omega_{a}$ and linewidth $\kappa_{a}$. In the dispersive regime the interaction takes the longitudinal form
$\hat{H}_{\mathrm{int}}=\chi_{a}\,\hat{a}^{\dagger}\hat{a}\,\hat{\sigma}_{z},$
so that the qubit transition frequency is shifted by $\pm\chi_{a}n$ conditional on the photon number $n$ (cross-Kerr shift $\chi_{a}$).
Photon-number fluctuations of the thermal mode are modeled as stationary Gaussian noise with exponential correlations,
$C_{nn}(\tau)\equiv\big\langle\delta\hat{n}(\tau)\,\delta\hat{n}(0)\big\rangle=\bar{n}(\bar{n}+1)\,e^{-\kappa_{a}|\tau|},\qquad\delta\hat{n}\equiv\hat{a}^{\dagger}\hat{a}-\bar{n},$
where $\bar{n}=[\exp(\hbar\omega_{a}/k_{B}T)-1]^{-1}$. Equivalently, the symmetrized noise spectrum is
$S_{nn}(\omega)=\int_{-\infty}^{\infty}d\tau\,e^{i\omega\tau}\,C_{nn}(\tau)=\frac{2\kappa_{a}\,\bar{n}(\bar{n}+1)}{\kappa_{a}^{2}+\omega^{2}}.$ (31)
During a Ramsey sequence of duration $\tau$, the qubit accumulates a stochastic phase
$\phi(\tau)=2\chi_{a}\int_{0}^{\tau}\!dt\,\hat{n}(t)=2\chi_{a}\bar{n}\,\tau+2\chi_{a}\int_{0}^{\tau}\!dt\,\delta\hat{n}(t).$ (32)
Averaging over Gaussian fluctuations using the cumulant expansion, the off-diagonal coherence is
$C_{q}(\tau)=\frac{\rho_{eg}(\tau)}{\rho_{eg}(0)}=e^{-i2\chi_{a}\bar{n}\tau}\;\exp\!\left[-\tfrac{1}{2}\big\langle\!\big\langle\phi_{\mathrm{noise}}^{2}(\tau)\big\rangle\!\big\rangle\right],$
with
$\big\langle\!\big\langle\phi_{\mathrm{noise}}^{2}(\tau)\big\rangle\!\big\rangle=(2\chi_{a})^{2}\int_{0}^{\tau}\!\!\int_{0}^{\tau}dt_{1}\,dt_{2}\,C_{nn}(t_{1}-t_{2}).$
Carrying out the double integral yields
$\int_{0}^{\tau}\!\!\int_{0}^{\tau}\!dt_{1}\,dt_{2}\,e^{-\kappa_{a}|t_{1}-t_{2}|}=2\int_{0}^{\tau}\!d\theta\,(\tau-\theta)\,e^{-\kappa_{a}\theta}=\frac{2\big(\kappa_{a}\tau-1+e^{-\kappa_{a}\tau}\big)}{\kappa_{a}^{2}}.$
The Ramsey coherence envelope is therefore
$C_{q}(\tau)=\exp\!\Big[-i\,2\chi_{a}\bar{n}\,\tau\Big]\;\exp\!\Big[-(2\chi_{a})^{2}\,\bar{n}(\bar{n}+1)\,f_{\kappa}(\tau)\Big],$ (33)
with
$f_{\kappa}(\tau)=\frac{\kappa_{a}\tau-1+e^{-\kappa_{a}\tau}}{\kappa_{a}^{2}}.$ (34)
This expression interpolates smoothly between $f_{\kappa}(\tau)\!\simeq\!\tfrac{1}{2}\tau^{2}$ for $\kappa_{a}\tau\!\ll\!1$ (quasi-static diffusion) and $f_{\kappa}(\tau)\!\simeq\!\tau/\kappa_{a}$ for $\kappa_{a}\tau\!\gg\!1$ (Markovian exponential decay) [35, 33, 1].
From Eq. (33), the amplitude is $C_{q}(\tau,T)=e^{-\Gamma_{q}(T)}$, with
$\Gamma_{q}(T)=(2\chi_{a})^{2}\bar{n}(\bar{n}+1)f_{\kappa}(\tau),$
and the phase is $\Phi_{q}(\tau,T)=2\chi_{a}\tau\bar{n}$. The derivatives are
$\partial_{T}C_{q}=-\,e^{-\Gamma_{q}}\,\Gamma_{q}^{\prime}(T),\qquad\Gamma_{q}^{\prime}(T)=(2\chi_{a})^{2}f_{\kappa}(\tau)\,\big(1+2\bar{n}(T)\big)\,\partial_{T}\bar{n}(T),$
and
$\partial_{T}\Phi_{q}(\tau,T)=2\chi_{a}\,\tau\,\partial_{T}\bar{n}(T).$ (35)
Substituting into Eq. (13), the QFI becomes
$\mathcal{F}_{q}=\frac{(\partial_{T}\Gamma_{q})^{2}}{e^{2\Gamma_{q}}-1}+e^{-2\Gamma_{q}}\big(2\chi_{a}\tau\big)^{2}\,\big(\partial_{T}\bar{n}\big)^{2}.$ (36)
[…]
Importantly, in this protocol the qubit itself is the sensor: it accumulates the thermal information through the same dispersive coupling that simultaneously dephases it. This trade-off is intrinsic to the qubit-only scheme and cannot be eliminated by echo or refocusing, since the stochastic diffusion encodes the very information being extracted. Thus, practical operation requires choosing $\tau\approx\tau^{\star}$ and balancing the gain from large $\chi_{a}$ against the inevitable loss of visibility. In contrast, the coherence-mediated scheme off-loads the sensing burden onto a high-$Q$ probe mode: while the probe collects the temperature-dependent phase diffusion, the qubit can be dynamically decoupled from it, enabling much longer interaction times and leveraging the stability of a 3D cavity to preserve visibility.
Appendix E Sixth–order perturbation theory for the two–qubit mediated cross–Kerr
[…]
$H_{0}/\hbar=\omega_{a}a^{\dagger}a+\omega_{b}b^{\dagger}b+\tfrac{\omega_{1}}{2}\sigma_{z1}+\tfrac{\omega_{2}}{2}\sigma_{z2},$
and the interaction term
$V/\hbar=g_{a1}(a\sigma_{1}^{+}+a^{\dagger}\sigma_{1}^{-})+g_{b2}(b\sigma_{2}^{+}+b^{\dagger}\sigma_{2}^{-})+J_{XY}(\sigma_{1}^{+}\sigma_{2}^{-}+\sigma_{1}^{-}\sigma_{2}^{+}).$
In the dispersive regime, $|\Delta_{a1,b2}|\gg|g_{a1,b2}|,\text{and}|\Delta_{12}|\gg|J_{XY}|$, with detunings $\Delta_{a1}=\omega_{a}-\omega_{1},\Delta_{b2}=\omega_{b}-\omega_{2},\Delta_{12}=\omega_{1}-\omega_{2}.$
The connected 6th-order perturbation theory for the energy correction can be calculated as:
[…]
  1. 1.
[…]
  2. 2.
[…]
  3. 3.
[…]
  4. 4.
[…]
  5. 5.
[…]
  6. 6.
[…]
Collecting these steps, the representative ladder can be summarized as:
[…]
The product of the matrix elements (numerator) for the path is straightforward
[…]
Each intermediate state contributes an energy denominator $E_{i}-E_{j,k,l,m,n}$:
[…]
and the denominator product is
$\mathcal{D}=\Delta_{a1}^{2}\,(\Delta_{a1}+\Delta_{12})^{2}\,(\Delta_{a1}+\Delta_{b2}).$
The derivation above follows the ladder where mode $a$ excites $Q_{1}$ first. There exists a mirror ladder where mode $b$ excites $Q_{2}$ first, producing an analogous expression with indices $a_{1}\leftrightarrow b_{2}$. Therefor, the 6th–order correction contributes to an effective cross–Kerr interaction
$H_{\mathrm{eff}}\supset\lambda\,a^{\dagger}a\,b^{\dagger}b,\qquad\lambda\;\sim\;\frac{g_{a1}^{2}g_{b2}^{2}J_{XY}^{2}}{\Delta_{a1}^{2}(\Delta_{a1}+\Delta_{12})^{2}(\Delta_{a1}+\Delta_{b2})}\;+\;\text{(mirror and permutations)}.$
Additionally, there are other time orderings permitted by Rayleigh–Schrödinger perturbation theory. Summing all allowed ladders restores a symmetric expression in $(\Delta_{a1},\Delta_{b2},\Delta_{12})$, and gives the effective cross Kerr
$\lambda=\frac{8\chi_{a1}\chi_{b2}J_{XY}^{2}}{\Delta_{12}^{3}},$ (38)
with $\chi_{a1}=-g_{a1}^{2}/\Delta_{a1}$ and $\chi_{b2}=-g_{b2}^{2}/\Delta_{b2}$.
Appendix F Parasitic Dispersive Couplings and Qubit Dephasing
While the cross-Kerr interaction between the probe ($b$) and thermal ($a$) modes is the desired sensing channel, the qubit also has residual dispersive couplings to each mode, with strengths $\chi_{b}$ and $\chi_{a}$. These couplings do not encode thermometric information; they act as parasitic dephasing pathways that reduce Ramsey visibility during the readout window of duration $\tau_{R}$.
If the qubit couples dispersively to a bosonic mode $m$ with photon-number variance $\mathrm{Var}(n_{m})$ and linewidth $\kappa_{m}$, a second-order cumulant expansion gives the coherence envelope during the Ramsey window [35]
$C_{q}^{(m)}(\tau_{R})=\exp\!\left\{-\,4\chi_{m}^{2}\,\mathrm{Var}(n_{m})\left[\frac{\kappa_{m}\tau_{R}-1+e^{-\kappa_{m}\tau_{R}}}{\kappa_{m}^{2}}\right]\right\}.$ (39)
Equation (39) smoothly interpolates between Gaussian decay in the quasi-static limit ($\kappa_{m}\tau_{R}\ll 1$) and exponential decay in the Markovian limit ($\kappa_{m}\tau_{R}\gg 1$).
For the high-$Q$ 3D cavity probe, $\kappa_{b}\tau_{R}\ll 1$, so (39) reduces to
$C_{q}^{(b)}(\tau_{R})\approx\exp\!\left[-2\chi_{b}^{2}\,\mathrm{Var}(n_{b})\,\tau_{R}^{2}\right],$ (40)
with photon-number variance for a displaced thermal state
$\mathrm{Var}(n_{b})=\bar{n}_{b}(1+\bar{n}_{b})+|\alpha|^{2}(1+2\bar{n}_{b}).$
Here $|\alpha|^{2}$ is the coherent population and $\bar{n}_{b}$ the residual thermal occupation. Increasing $|\alpha|$ boosts temperature-to-phase transduction in the probe but also increases parasitic dephasing during the Ramsey readout via $\chi_{b}$. The information encoded in the probe during the imprinting interval is independent of $\chi_{b}$; $\chi_{b}$ primarily sets the speed and backaction of the readout.
Qubit dephasing due to residual coupling to the thermal mode also follows Eq. (39), with
$\mathrm{Var}(n_{a})=\bar{n}_{a}(1+\bar{n}_{a}).$
In practice the thermal mode is implemented as a low-$Q$ resonator with $\kappa_{a}\tau_{R}\gtrsim 1$, placing Eq. (39) in the Markovian regime. Expanding for $\kappa_{a}\tau_{R}\gg 1$ yields
$C_{q}^{(a)}(\tau_{R})\approx\exp\!\left[-\,\frac{4\chi_{a}^{2}\,\mathrm{Var}(n_{a})}{\kappa_{a}}\,\tau_{R}\right],$ (41)
i.e., an exponential decay with rate $\gamma_{\varphi}^{(a)}=4\chi_{a}^{2}\,\mathrm{Var}(n_{a})/\kappa_{a}$. Thus, a strongly damped thermal resonator acts as a broadband Markovian bath: while $\mathrm{Var}(n_{a})$ grows with temperature, the large linewidth $\kappa_{a}$ suppresses the overall dephasing rate. This contrasts with the high-$Q$ limit, where quasi-static fluctuations induce Gaussian-in-time visibility loss.
In summary, both $\chi_{b}$ and $\chi_{a}$ enter through the general law (39), but with distinct photon statistics and dynamical regimes. Engineering small residual couplings ensures that visibility is limited only during the short Ramsey window, while the probe amplitude $\alpha$ can be used as a controlled resource to optimize sensitivity without incurring unnecessary readout dephasing.
Appendix G Phase-to-Quadrature Mapping
Figure 8: Phase–to–quadrature mapping pulse sequence. A short pulse prepares the high-$Q$ 3D cavity (probe) in a coherent state of amplitude $\alpha$ (blue). The probe then interacts with the thermal mode $a$ for a duration $\tau$, acquiring a phase $\phi_{b}=\lambda\!\int_{0}^{\tau}n_{a}\,dt$ (yellow), while the qubit remains idle. Next, a short sensing window of length $t_{R}$ turns on the qubit–probe dispersive coupling $\chi_{b}$ (green). During this window a weak LO displacement $\beta=|\beta|e^{i\theta}$ is applied to set the analyzed quadrature (green). A Hahn-echo Ramsey block ($\pi/2\;–\;\pi\;–\;\pi/2$) inside the window, with a $\pi$ phase flip of the LO at the echo, cancels phase-blind terms and maps the probe quadrature $\hat{X}_{\theta}$ onto a small qubit $Z$-phase (red).
In the coherence-mediated scheme, the coherent probe accumulates a phase shift $\phi_{b}=\lambda\!\int_{0}^{\tau}n_{a}\,dt$ during its interaction with the thermal mode over the interaction time $\tau$. However, its photon number remains fixed, $n_{b}=|\alpha e^{i\phi_{b}}|^{2}=|\alpha|^{2}$, so a plain Ramsey evolution under the dispersive coupling $\chi_{b}\hat{n}_{b}\hat{\sigma}_{z}$ is phase blind.
To read out the probe’s phase fluctuations, we adopt a short, phase-sensitive mapping block [57, 58]. The thermal mode $a$ imprints a random phase $\phi_{b}$ on the probe $b$ during $\tau$; a following brief mapping then converts the probe’s quadrature $X_{\theta}$ into a small qubit $Z$-phase. Shot-to-shot fluctuations of $\phi_{b}$ translate into fluctuations of the qubit phase and reduce the Ramsey visibility. From this visibility we infer $\mathrm{Var}(\phi_{b})$, then $\bar{n}_{a}(T)$, and finally the temperature $T$.
During the short sensing window of duration $t_{s}$ between the two $\pi/2$ pulses, we apply a tiny local-oscillator (LO) displacement on the probe, $\beta=|\beta|e^{i\theta}$. The photon number seen by the qubit becomes
$n_{b}=\big|\alpha e^{i\phi_{b}}+\beta e^{i\theta}\big|^{2}=|\alpha|^{2}+|\beta|^{2}+2|\alpha||\beta|\cos(\phi_{b}-\theta),$
which contains a phase-sensitive cross-term. With a Hahn echo inside the window and a $\pi$ flip of the LO phase ($\theta\!\to\!\theta+\pi$) at the echo, the phase-blind (number-like) contributions cancel while the quadrature term adds. The qubit therefore acquires a small, phase-sensitive angle
$\theta_{q}(\phi_{b})=g_{m}\,2|\alpha|\cos(\phi_{b}-\theta),\qquad g_{m}=2|\beta|\,\chi_{b}\,\tau_{s}.$
The net effect is an effective, brief longitudinal interaction that maps the probe quadrature onto a small qubit $Z$-phase. The mapping strength is set by a single, experimentally convenient gain knob $g_{m}$, and the analyzed quadrature is set by the LO phase $\theta$ (typically biased near the maximum-slope point, $\phi_{b}-\theta\!\approx\!\pi/2$).
Operationally, the sequence is: (i) let the probe interact with the thermal mode for a time $\tau$ while the qubit is idle, so the probe acquires a random phase $\phi_{b}(T)$; (ii) turn on the mapping window $t_{R}$ bracketed by $\pi/2$ pulses, with the qubit–probe dispersive coupling active and a weak LO on the probe; (iii) insert a Hahn echo on the qubit and flip the LO phase by $\pi$ midway through the window. The echo removes phase-blind terms, while the LO phase flip preserves the desired quadrature term. The qubit then accumulates a tiny, phase-sensitive angle proportional to the probe quadrature at phase $\theta$. Fig. 8 illustrates a pulse sequence for this phase-to-quadrature mapping scheme.
Two practical remarks follow. First, the mapping is kept in the “small-angle” regime: the LO injects $\ll 1$ photon on average and the window is much shorter than $T_{2}$, so backaction and higher-order corrections are negligible. Second, the same idea can be realized without an explicit LO by parametrically modulating the qubit (or a tunable coupler) at the probe frequency to generate a direct longitudinal coupling to $X_{\theta}$ during $t_{s}$ [52, 31].